Calculate n (in 2008) for carbon-14 in charred plant remains for two different eruptions of Mt. Vesuvius, 472 AD and 512 AD. (ti/2 for 14 5730 y) Number 472 AD, n Number 512 AD, n
Answer #1

The half life equation

A = A0*(1/2)^(t/L)

n = A/A0

n = (0.5)^(t/L)

at

t = 472 AD --> (2008-472) = 1536

n = (0.5)^(t/L)

n = (0.5)^(1536/5730)

n = 0.83

now...

t = 512 AD --> (2008-512 ) = 1496

n = (0.5)^(t/L)

n = (0.5)^(1496/5730)

n = 0.83446